# Frontiers in Probabilistic Inference: Sampling Meets Learning

### v1

@@ Abstract
Simulating the conditioned dynamics of diffusion processes, given their initial and terminal states, is an important but challenging problem in the sciences. The difficulty is particularly pronounced for rare events, for which the unconditioned dynamics rarely reach the terminal state. In this work, we leverage a self-consistency property of the conditioned dynamics to learn the diffusion bridge in an iterative online manner, and demonstrate promising empirical results in a range of settings.
@@ 1 Introduction
Diffusion processes are ubiquitous throughout mathematics, science, and, more recently, machine learning. A diffusion process $(X_{t})_{t\in[0,T]}$ in ${\mathbb{R}}^{d}$ is governed by a stochastic differential equation
for drift $b:[0,T]\times{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ and diffusion coefficient $\sigma:[0,T]\times{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d\times d}$. In this work we will consider elliptic diffusions, so $\sigma\sigma^{\top}$ is of full-rank, and make the standard assumptions that $b,\sigma$ are Lipschitz continuous and of linear growth.
Given the SDE dynamics (1), one can use time-discretisation schemes such as the Euler-Maruyama method to simulate samples of the diffusion process. However, often in applications it is required to simulate the conditioned dynamics of this underlying reference process, conditional on terminating at a point $X_{T}=x_{T}$. The sampling of such diffusion bridges has applications in chemistry [7, 17], finance [19, 16], and shape evolution [1] to name a few. This is typically a challenging problem, particularly when the conditioned event in question occurs rarely under the base dynamics (1).
@@
Traditional approaches for sampling diffusion bridges have commonly used MCMC or SMC-based methods [3, 21]. More recently, neural approximations have been used to learn the control drift of the conditioned dynamics, leveraging tools from score-matching [23, 2] or Malliavin calculus [35]. However, these methods learn the neural approximation using samples from uncontrolled dynamics, limiting their success for rare conditioning events as the training samples may not adequately cover the support of the diffusion bridge solution. Motivated by this, [45] learn a neurally-guided bridge by minimising the KL-divergence to the true conditioned dynamics, though they require backpropagation through the simulated trajectories which can impact scalability.
@@
In this work, we derive a self-consistency property of the diffusion bridge solution in terms of the control drift $u(t,X_{t})$ and the Jacobian process $J_{t|s}$ along the solution trajectories. We consider a class of neural-controlled diffusion processes that ensures transportation to the conditioned endpoint $x_{T}$, and train according to a proposed family of self-consistency losses based on the above property of the solution. These fixed-point-style losses are trained in an online manner, and provide scalable training without the need to backpropagate through the simulated trajectories. We demonstrate the validity of our proposed approach in a range of settings, showing competitive performance with existing methods at a lower computational cost.
Figure 1: The self-consistency property $u(s,X_{s})=\mathbb{E}_{\mathbb{P}^{u}}[u(t,X_{t})J_{t|s}|X_{s}]$ relates the control value at an earlier time $s$ to its value at a later time $t$, along the trajectories.
Figure 2: The diffusion bridge is the unique process of form (3) bridging from $x_{0}$ to $x_{T}$ with $u$ of gradient form, that satisfies the self-consistency property (6).
@@ 2 Background
Let us denote by $\mathbb{P}$ the path measure induced by the unconditioned reference process $X_{t}$ in (1), and consider a ‘target’ path measure $\mathbb{Q}$} induced by a reweighting density function $F(X_{T})$ defined on the terminal state,
@@
It is a well-known result from stochastic analysis [18, Section 1.3] that samples from $\mathbb{Q}$} can be obtained by simulating the controlled diffusion process $X_{t}^{u}$, whose dynamics are governed by the SDE
${\mathrm{d}}X_{t}^{u}=\big[b(t,X_{t}^{u})+(\sigma\sigma^{\top})(t,X_{t}^{u})u(t,X_{t}^{u})\big]{\mathrm{d}}t+\sigma(t,X_{t}^{u})\,{\mathrm{d}}B_{t},\qquad X_{0}^{u}=x_{0},$ (3)
where the control drift is given by Doob’s $h$-transform, $u(t,x)=\nabla_{x}\log{\mathbb{E}}_{\mathbb{P}}[F(X_{T})|X_{t}=x]$. Observe that diffusion bridges are recovered from this framework by using the singular reweighting function $F(X_{T})=\delta_{x_{T}}(X_{T})$, which intuitively corresponds to ‘retaining only trajectories that terminate at $x_{T}$’. In this case, Doob’s $h$-transform takes the more familiar form $u(t,x)=\nabla_{x}\log p(T,x_{T}|t,x)$, where $p(t,y|s,x)$ denotes the transition densities of the base process $X_{t}$.
@@
In order to explain our approach, we now introduce the Jacobian, or derivative, process $J_{t|s}=\nabla_{X_{s}}X_{t}$ [39, Chapter V.13], which evolves for $t>s$ according to the SDE
${\mathrm{d}}J_{t|s}=\nabla b(t,X_{t})J_{t|s}\,{\mathrm{d}}t+\nabla\sigma(t,X_{t})J_{t|s}\,{\mathrm{d}}B_{t},\qquad J_{s|s}=\mathrm{Id}.$ (4)

### v2

# Control Consistency Losses for Diffusion Bridges

@@ Abstract
Simulating the conditioned dynamics of diffusion processes, given their initial and terminal states, is an important but challenging problem in the sciences. The difficulty is particularly pronounced for rare events, for which the unconditioned dynamics rarely reach the terminal state. In this work, we propose a novel approach for learning diffusion bridges based on a self-consistency property of the optimal control. The resulting algorithm learns the conditioned dynamics in an iterative online manner, and exhibits strong performance in a range of empirical settings without requiring differentiation through simulated trajectories. Beyond the diffusion bridge setting, we draw connections between our self-consistency framework and recent advances in the wider stochastic optimal control literature.
@@ 1 Introduction
Diffusion Processes
Diffusion processes are ubiquitous throughout mathematics, science, and, more recently, machine learning. A diffusion process $(X_{t})_{t\in[0,T]}$ in ${\mathbb{R}}^{d}$ is governed by a stochastic differential equation (SDE),
for a drift coefficient $b:[0,T]\times{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ and diffusion coefficient $\sigma:[0,T]\times{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d\times d}$. In this work we will consider elliptic diffusions, so $\sigma\sigma^{\top}$ is of full-rank, and make the standard assumptions that $b,\sigma$ are Lipschitz continuous and of linear growth.
@@
Given the SDE dynamics (1), one can use standard time-discretisation schemes such as the Euler–Maruyama method to simulate samples of the unconditioned diffusion process $X_{t}$. However, often in applications one is instead interested in simulating the conditioned dynamics of this underlying base process, conditional on terminating at a known location $X_{T}=x_{T}$ (Baudoin, 2002). Such diffusion bridges play an important role in a wide range of domains, including chemistry (Bolhuis et al., 2002; E & Vanden-Eijnden, 2010), finance (Elerian et al., 2001; Durham & Gallant, 2002), and shape evolution (Arnaudon et al., 2022) to name a few. Sampling diffusion bridges is typically challenging, particularly when the conditioned event in question occurs rarely under the base dynamics (1).
@@
Traditional approaches for diffusion bridge sampling have commonly used MCMC or SMC-based methods (Beskos et al., 2008; Fearnhead et al., 2008). More recently, neural approximations have been used to learn the controlled drift of the conditioned dynamics, leveraging tools from score-matching (Heng et al., 2025; Baker et al., 2025) or Malliavin calculus (Pidstrigach et al., 2025). Once trained, such neural approaches enable direct simulation of samples from the diffusion bridge. However, these methods learn the neural approximation using samples from uncontrolled dynamics, limiting their success for rare conditioning events as the training samples may not adequately cover the region near the terminal state $x_{T}$. Motivated by this, Yang et al. (2025) learn a neurally-guided bridge by minimising the KL-divergence to the true conditioned dynamics, though they require backpropagation through the simulated trajectories which can impact scalability.
@@
In this work, we derive a self-consistency property of the diffusion bridge solution in terms of the control drift $u(t,X_{t})$ and the Jacobian process $J_{t|s}$ along the solution trajectories (Theorem 2.1, Figure 2). In combination with a terminal condition that enforces bridging to the known endpoint $x_{T}$ (Theorem 3.1, Figure 2), we train a neural-controlled diffusion process using a proposed family of self-consistency losses based on the above property of the solution. These fixed-point-style losses are trained in an online manner, and provide scalable training without the need to backpropagate through the simulated trajectories. We demonstrate strong empirical performance in a range of settings, competitive with state-of-the-art existing methods at a significantly reduced computational cost. Additionally, we explain how our self-consistency framework extends to wider stochastic optimal control problems, offering new perspectives and connections to recent developments.
Figure 1: Theorem 2.1: The self-consistency property (SC1) relates the control value at an earlier time $s$ to its value at a later time $t$, along the solution trajectories.
Figure 2: Theorem 3.1: The diffusion bridge is the unique process of form (3) bridging from $x_{0}$ to $x_{T}$ with $u$ of gradient form, that satisfies the self-consistency property (SC1).
@@ 2 Background
We begin with a more general setting, from which diffusion bridges will be recovered as a limiting case. Let us denote by $\mathbb{P}$ the path measure associated with the unconditioned reference process $X_{t}$ in (1), and consider a ‘target’ path measure $\mathbb{Q}$} induced by a reweighting density function $F(X_{T})$ defined on the terminal state,
@@
It is a well-known result from stochastic analysis (Eberle, 2019, Section 1.3) that samples from $\mathbb{Q}$} can be obtained by simulating a controlled diffusion process $X_{t}^{u}$, with dynamics given by the following SDE started at $X_{0}^{u}=x_{0}$,
${\mathrm{d}}X_{t}^{u}$ $=\big(b+(\sigma\sigma^{\top})u\big)(t,X_{t}^{u})\,{\mathrm{d}}t+\sigma(t,X_{t}^{u})\,{\mathrm{d}}B_{t}.$ (3)
Moreover, the optimal control [^1] Henceforth, we will use the term optimal control to refer to the $u$ that implements the conditioning $X_{T}=x_{T}$; the connection to optimal control more broadly will be explained in Appendix F. $u^{*}$ (that is, the control achieving (2)), is given by Doob’s $h$-transform,
$u^{*}(t,x)=\nabla_{x}\log{\mathbb{E}}_{\mathbb{P}}[F(X_{T})|X_{t}=x].$ (4)
Observe that diffusion bridges are recovered from this framework by using the singular reweighting function $F(X_{T})=\delta_{x_{T}}(X_{T})$}, which intuitively corresponds to ‘retaining only trajectories that terminate at $x_{T}$’. Plugging this into (4), Doob’s $h$-transform takes the more familiar form $u^{*}(t,x)=\nabla\log p(T,x_{T}|t,x)$, where $p(t,y|s,x)$ denotes the transition densities of the base process $X_{t}$.
@@
In order to explain our approach, we now introduce the Jacobian, or derivative, process $J_{t|s}=\nabla_{X_{s}}X_{t}$ (Rogers & Williams, 2000, Chapter V.13), which evolves for $t>s$ according to the SDE
${\mathrm{d}}J_{t|s}=\nabla b(t,X_{t})J_{t|s}{\mathrm{d}}t+\nabla\sigma(t,X_{t})J_{t|s}{\mathrm{d}}B_{t},~~J_{s|s}=Id.$ (5)